maintnance period progress

From: Flemming Videbaek (videbaek@sgs1.hirg.bnl.goV)
Date: Wed Oct 10 2001 - 21:03:31 EDT

  • Next message: Pawel Staszel: "DC settings"

    I will like to give you a briefing on the work/activities over the 48hour period with no RHIC beam- the list is not in
    any prioritized order.
    
    - L4032 HV system repairs. Bob Scheetz identified a partial short (39Ohm) within a cable connecting MF 2 & 3- hopefully
    this is the culprit that has caused so many problemsr.
    - the FFS MF 1440 comm. was re-established recycling the power. (RAS)
    -The bad H2 channels was traced by CEJ/DO. Some additional problems with the BFS 1440 and a bad switch.
    -Exchange of some bad FEE cards by JIJ.In progress and evaluation- will likely move row 19/20 FEE cards on TPM2
    to row 17/18 to get a better response in the drift direction.
    - Two panels with 5 horizontal; fibre each has been installed behind T1 and in front of T2.(DB and E.baker). The DAQ and HV was modified (FV). This project is now complete.
    - The HV distribution for T5 is beeing redone to minimize module dependence in case of failure and reduce curtent load on each HV channel.
    - The TPM1 Drift Velocity Monitor(DVM1) is moved to behind D5 to avoid exposure to the beam scattered particles.
      The T1 that was 'fried' is being removed for replacement.
    The area should be swept tomorrow eraly evening, and a checkout round should be made before then.
    
    
    
    
    ------------------------------------------------------
    Flemming Videbaek
    Physics Department
    Brookhaven National Laboratory
    
    tlf: 631-344-4106
    fax 631-344-1334
    e-mail: videbaek@bnl.gov
    



    This archive was generated by hypermail 2b30 : Wed Oct 10 2001 - 21:02:26 EDT